# Mechanical memories in solids, from disorder to design

### v1

@@ 4.3 Coupled hysterons in mechanical metamaterials
The perfect return-point memory of Fig. 5d indicates that interactions between hysterons in this system are negligible [65]. This lack of coupling may be desirable for straightforward control of a system’s state, as in the Preisach model. But as with amorphous solids in Sec. 3, interactions make many more behaviors possible. Metamaterials are an opportunity to tailor these interactions and design mechanical memories and computations at the limits of glassy matter [76, 77, 78].
To demonstrate tunable interactions, Bense and van Hecke [74] made a corrugated silicone sheet in the shape of a cylindrical arc with three sinusoidal ridges (Fig. 6a). As the sheet is compressed and released, each ridge acts as a hysteron that buckles and straightens, causing the jumps in the three nested force versus displacement curves of Fig. 6b. Initially, just four states are reachable, shown in Fig. 6c with their transitions. However, in this design the switching thresholds—and hence the effects of interactions—can be tuned via the boundaries. After tilting the bottom plate by a small angle ($\mathcal{O}(10)$ milliradians), one obtains the dramatically different transition graph in Fig. 6d. More states are reachable, and an instance of “scrambling” is observed, shown by the double-lined blue arrows. Comparing the down transitions from the states 111 and 110, one may probe the influence of the third hysteron (in the 1 or 0 state, respectively) on the first two hysterons (which both begin as 1 in this pair of states). The effect is that the second hysteron switches first when the third hysteron is ‘on’ ($111\rightarrow 101$), but the first hysteron switches first when the third hysteron is ‘off’ ($110\rightarrow 010$), making return-point memory impossible. This is an example where interactions are strong enough that their presence can be inferred from a transition graph. Other metamaterials are also being designed with strong interactions between bistable hysteretic elements, with an eye towards performing logic operations in materia [79, 80].
@@ 4.5 Metamaterials that defy coupled-hysteron models
Finally, some metamaterials with bistable elements can push beyond the limits of hysteron models by bringing in other physical interactions. Buckled beams in a material can resemble hysterons [79], but work by Ding, Kwakernaak, and van Hecke [86, 87] describes systems where specially-designed contacts between beams are made, broken, and evolve under friction continuously in time. Under cyclic driving, these systems show discrete memory behaviors of the kind described above—but the kinematics and mechanisms of these behaviors are described poorly by discrete (hysteron-based) degrees of freedom.
@@ 5 MEMORY AND DEGREE OF DISORDER: FROM AMORPHOUS SOLIDS TO POLYCRYSTALS
Many non-designed materials feature large crystalline domains unlike anything found in an amorphous solid. Indeed, the conventional starting point for construction theories of solids has been to treat these materials as extensions of the tremendously successful theories of crystalline matter. Thus it is a priori unclear whether amorphous or crystalline matter is the better prototype for memory in polycrystals.
@@
Figure 7: Bulk mechanics and mesoscopic rearrangements in polycrystals. (a) Ratio of shear modulus to bulk modulus for five systems that span the spectrum from totally disordered ($F_{6}\approx 0$) to perfect crystal ($F_{6}=1$), where $F_{6}$ measures the correlations of bond-orientational order among neighboring particles. Four of the systems exhibit a vanishing ratio of shear to bulk modulus ($G/B$) in the limit of vanishing pressure, a hallmark of the jamming transition in amorphous solids. Only the system with $F=0.95$ exhibits the plateau in $G/B$ that is characteristic of a crystal. (b) Comparison of rearranging regions in a more amorphous sample (left) and a more polycrystalline sample (right). Particles are colored according to the degree of hysteresis in their response to shear (thresholds $\gamma^{+}$ and $\gamma^{-}$ are labeled $\gamma_{\text{on}}$ and $\gamma_{\text{off}}$). The appearances and statistics of rearrangements are similar between sample types. (c) Two-point cluster function from these two systems, showing how the average correlation within a rearranging region decays with distance from the region’s center. The data from the amorphous solid and from the polycrystal track each other remarkably well. Panel a courtesy of Carl Goodrich [88]. Panels b and c reproduced with permission from Ref. 89.
@@
To elucidate this issue, Goodrich et al. [88] produced a set of three-dimensional jammed systems along a spectrum from completely disordered to crystalline. They found that adding only a small amount of disorder to a crystal brings its behaviors closely in line with the disordered case (Fig. 7a), consistent with other studies [90, 91]. This view of polycrystals extends to their plastic response: in simulations they exhibit localized rearrangements under loading [90, 92, 93], which arise in comparatively softer regions [94]. In experiments, Keim and Arratia [89] created 2D jammed solids similar to the amorphous samples in Fig. 2, but switched from bidisperse particle sizes to monodisperse, forming crystalline regions separated by disordered grain boundaries. Remarkably, localized rearrangements are nearly the same diameter and magnitude whether the sample is amorphous or polycrystalline (Fig. 7b,c); the main difference is simply that the polycrystal has fewer of them. These rearrangements are predominantly at grain boundaries [89, 95], although particles within crystalline regions can rearrange as well [33]. These microscopic results suggest that polycrystals can form memories of cyclic driving. Indeed, this is in line with experiments showing memories in a bubble raft in an annular geometry, where one can see crystalline domains [14]. How the strength of the memory signal, or the capacity for multiple memories, varies with the degree of disorder in polycrystals remain open avenues of study. For instance, amorphous solids appear to present strong memories of cyclic driving near yielding [14], whereas 2D Lennard-Jones polycrystals can coarsen into monocrystals as yielding is approached [96, 97]—raising the possibility that near yielding, polycrystals can lose their memory capacity altogether.
@@
Intriguingly, while the crystalline domains in a polycrystal seem to have few of the rearrangements that might encode memories of amplitude, they do store at least one other kind of memory: the direction of the last large shear when a sample was prepared. Teich et al. [33] analyze experimental data from 2D jammed amorphous bidisperse particle systems [17, 89]. They measure slight anisotropies at the particle level, similar to the amorphous solid results in Fig. 3(a, b)—but here, the memory of direction is strongest within crystalline domains, where it corresponds to a slight distortion of the lattice [33]. This result shows that crystallinity and mechanical memory are not mutually exclusive.
@@ 6.1 Dynamics
Particle rearrangements are not instantaneous, due to inertial, viscous, or thermal considerations [46, 98, 99, 100]—nor are these timescales uniform in a disordered system. Away from the quasistatic limit, some relaxations may not keep up with a deformation. The influence of a “slow” rearrangement on its neighbors will change when the strain rate is increased, or if a turning point of a cyclic deformation is reached too soon [101, 102], so that when the material comes to rest it encodes past dynamics. Like extended transients and multiperiodicity, this sensitivity may even play out over many cycles [102]. One exciting possibility is that a suitable readout protocol could recover the rate(s) of past deformations, a task that today is usually relegated to failure analysis [99, 47]. Some of these ideas may also apply to mechanical metamaterials, where it was recently shown that driving the material dynamically can allow one to reach states that would be unattainable using quasistatic deformation [103].
@@ ACKNOWLEDGMENTS
We are grateful to the many people who have shaped our understanding and appreciation of memory over years of interactions, several of whom also provided feedback on a draft of this article. A partial list must include Omri Barak, Karen Daniels, Mathew Diamond, Larry Galloway, Varda Hagh, Yoav Lahini, Chloe Lindeman, Craig Maloney, Alan Middleton, Muhittin Mungan, Sidney Nagel, Abigail Plummer, Ido Regev, Srikanth Sastry, Martin van Hecke, Damien Vandembroucq, Zhicheng Wang, Zorana Zeravcic, and organizers and participants of the May 2023 CECAM workshop “Pathways, Memory, and Emergent Computation in Nonequilibrium Systems,” Sde Boqer, Israel. NCK was supported in part by a Research Grant from the Human Frontier Science Program (Ref.-No: RGP0017/2021).

### v2

@@ 4.3 Coupled hysterons in mechanical metamaterials
The simple hierarchical structure of Fig. 5d suggests that interactions between hysterons in this system are negligible [69]. This state of affairs may be desirable for straightforward control of a system’s state, as in the Preisach model. But as with amorphous solids in Sec. 3, interactions make many more behaviors possible. Metamaterials are an opportunity to tailor these interactions and design mechanical memories and computations at the limits of glassy matter [90, 91, 92].
To demonstrate tunable interactions, Bense and van Hecke [88] made a corrugated silicone sheet in the shape of a cylindrical arc with three sinusoidal ridges (Fig. 6a). As the sheet is compressed and released, each ridge acts as a hysteron that buckles and straightens, causing the jumps in the three nested force-versus-displacement curves of Fig. 6b. Initially, just four states are reachable, shown in Fig. 6c with their transitions. However, in this design the switching thresholds—and hence the effects of interactions—can be tuned via the boundaries. After tilting the bottom plate by a small angle ($\mathcal{O}(10)$ milliradians), one obtains the dramatically different transition graph in Fig. 6d. More states are reachable, and an instance of “scrambling” is observed, shown by the double-lined blue arrows. Comparing the down transitions from the states 111 and 110, one may probe the influence of the third hysteron (in the 1 or 0 state, respectively) on the first two hysterons (which both begin as 1 in this pair of states). The effect is that the second hysteron switches first when the third hysteron is ‘on’ ($111\rightarrow 101$), but the first hysteron switches first when the third hysteron is ‘off’ ($110\rightarrow 010$), making return-point memory impossible. This is an example where interactions are strong enough that their presence can be inferred from a transition graph. Other material systems are being designed with strong interactions between bistable hysteretic elements [93], often with an eye towards performing logic operations in materia [94, 95]. One active avenue is the development of architectures that allow one to rationally target particular transition pathways [96].
@@ 4.5 Metamaterials that defy coupled-hysteron models
Finally, some metamaterials with bistable elements can push beyond the limits of hysteron models by bringing in other physical interactions. Buckled beams in a material can resemble hysterons [94], but work by Ding, Kwakernaak, and van Hecke [102, 103] describes systems where specially-designed contacts between beams are made, broken, and evolve under friction continuously in time. Although these systems show discrete memory behaviors of the kind described above, the kinematics and mechanisms of these behaviors are described poorly by discrete (hysteron-based) degrees of freedom.
@@ 5 MEMORY AND DEGREE OF DISORDER: FROM AMORPHOUS SOLIDS TO POLYCRYSTALS
Many non-designed materials feature large crystalline domains unlike anything found in an amorphous solid. Indeed, the conventional starting point for constructing theories of solids has been to treat these materials as extensions of the tremendously successful theories of crystalline matter. Thus it is a priori unclear whether amorphous or crystalline matter is the better prototype for memory in polycrystals.
@@
Figure 7: Bulk mechanics and mesoscopic rearrangements in polycrystals. (a) Ratio of shear modulus to bulk modulus for five systems that span the spectrum from totally disordered ($F_{6}\approx 0$) to perfect crystal ($F_{6}=1$), where $F_{6}$ measures the correlations of bond-orientational order among neighboring particles. Four of the systems exhibit a vanishing ratio of shear to bulk modulus ($G/B$) in the limit of vanishing pressure, a hallmark of the jamming transition in amorphous solids. Only the system with $F=0.95$ exhibits the plateau in $G/B$ that is characteristic of a crystal. (b) Comparison of rearranging regions in a more amorphous sample (left) and a more polycrystalline sample (right). Particles are colored according to the degree of hysteresis in their response to shear (thresholds $\gamma^{+}$ and $\gamma^{-}$ are labeled $\gamma_{\text{on}}$ and $\gamma_{\text{off}}$). The appearances and statistics of rearrangements are similar between sample types. (c) Two-point cluster function from these two systems, showing how the average correlation within a rearranging region decays with distance from the region’s center. The data from the amorphous solid and from the polycrystal track each other remarkably well. Panel a reproduced with permission from Ref. 104. Panels b and c reproduced with permission from Ref. 105.
@@
To elucidate this issue, Goodrich et al. [104] produced a set of three-dimensional jammed systems along a spectrum from disordered to crystalline. They found that adding only a small amount of disorder to a crystal brings its behaviors closely in line with the most disordered case (Fig. 7a), consistent with other studies [106, 107]. This view of polycrystals extends to their plastic response: in simulations they exhibit localized rearrangements under loading [106, 108, 109], which arise in comparatively softer regions [110]. In experiments, Keim and Arratia [105] created 2D jammed solids similar to the amorphous samples in Fig. 2, but switched from bidisperse particle sizes to monodisperse, forming crystalline regions separated by disordered grain boundaries. Remarkably, localized rearrangements are nearly the same diameter and magnitude whether the sample is amorphous or polycrystalline (Fig. 7b,c); the main difference is simply that the polycrystal has fewer of them. These rearrangements are predominantly at grain boundaries [105, 111], although particles within crystalline regions can rearrange as well [37]. These microscopic results suggest that polycrystals can form memories of cyclic driving. Indeed, this is in line with experiments showing memories in a bubble raft in an annular geometry, where one can see crystalline domains [14]. How the strength of the memory signal, or the capacity for multiple memories, varies with the degree of disorder in polycrystals remain open avenues of study. For instance, amorphous solids appear to present strong memories of cyclic driving near yielding [14], whereas 2D Lennard-Jones polycrystals can coarsen into monocrystals as yielding is approached [112, 113]—raising the possibility that near yielding, polycrystals can lose their memory capacity altogether.
@@
Intriguingly, while the crystalline domains in a polycrystal seem to have few of the rearrangements that might encode memories of amplitude, they do store at least one other kind of memory: the direction of the last large shear when a sample was prepared. Teich et al. [37] analyzed experimental data from 2D jammed amorphous bidisperse particle systems [21, 105]. They measured slight anisotropies at the particle level, similar to the amorphous solid results in Fig. 3(a, b)—but here, the memory of direction is strongest within crystalline domains, where it corresponds to a slight distortion of the lattice [37]. This result shows that crystallinity and mechanical memory are not mutually exclusive.
@@ 6.1 Dynamics
Particle rearrangements are not instantaneous, due to inertial, viscous, or thermal considerations [50, 114, 115, 116]—nor are these timescales uniform in a disordered system. Away from the quasistatic limit, some relaxations may not keep up with a deformation, leading to nonlinear wave propagation [117, 18] and to new possibilities for memory: the influence of a “slow” rearrangement on its neighbors will change when the strain rate is increased, or if a turning point of a cyclic deformation is reached too soon [118, 119], so that when the material comes to rest it encodes past dynamics. Like extended transients and multiperiodicity, this sensitivity may even play out over many cycles [119]. One exciting possibility is that a suitable readout protocol could recover the rates of past deformations, a task that today is usually relegated to failure analysis [115, 51]. Some of these ideas may also apply to mechanical metamaterials, where it was recently shown that driving the material dynamically can allow one to reach states that would be unattainable using quasistatic deformation [120].
@@ ACKNOWLEDGMENTS
We are grateful to the many people who have shaped our understanding and appreciation of memory over years of interactions, several of whom also provided feedback on a draft of this article. A partial list must include Omri Barak, Karen Daniels, Mathew Diamond, Larry Galloway, Varda Hagh, Yoav Lahini, Chloe Lindeman, Craig Maloney, Enzo Marinari, Alan Middleton, Muhittin Mungan, Sidney Nagel, Abigail Plummer, Ido Regev, Srikanth Sastry, Martin van Hecke, Damien Vandembroucq, Zhicheng Wang, Zorana Zeravcic, and organizers and participants of the May 2023 CECAM workshop “Pathways, Memory, and Emergent Computation in Nonequilibrium Systems,” Sde Boqer, Israel. NCK was supported in part by a Research Grant from the Human Frontier Science Program (Ref.-No: RGP0017/2021).